OpenStudy (anonymous):

square rot of 81 divided by square root 7????

OpenStudy (anonymous):

\[=(9\sqrt{7})/7\]

OpenStudy (anonymous):

polmak what does QED stand for?

OpenStudy (anonymous):

\[\sqrt{\frac{81}{\sqrt{7}}}=\frac{9}{7^{1/4}} \]

OpenStudy (radar):

\[(9\sqrt{7})/7\]